9、Let W and L denote the wage and the amount of labor employed, respectively. A firm faces the labor supply curve L = 2W - 6 and the marginal product of labor is given by MPL = 20 - L. The firm sells its output in a perfectly competitive market at $0.50 each.

(a) If this labor market is perfectly competitive, find the equilibrium employment, the equilibrium wage, and the number of unemployed people.

(b) Suppose that the government imposes a minimum wage of $7. Find the equilibrium employment and the number of unemployed people.

(c) If this labor market is controlled by a non-discriminating monopsonist, find the equilibrium employment, the equilibrium wage, and the number of unemployed people.

(d) If this labor market is controlled by a non-discriminating monopsonist and the government imposes a minimum wage of $7, find the equilibrium employment and the number of unemployed people.
Answer #1

a) is = 20-6 MPL=20-L, P=.5 a) Perfect compt VMR = woge (146)/2= 6, •5 (MPL)= 4+6)/2 05 (90-L= -5 L +3 10-05 L = 5Lt3 Ang Eqm20 DATE: 1 PAGE NO Monopsonist, then at Eqm VMPL = MCL MCL Marginal cost of Labor L+6)/2.L (05 2+ 3) = .5L²+36 M CL = ATCL |A